# Graham’s rearrangement conjecture beyond the rectification barrier

### v1

@@ 2. Notation and parameters
We use standard asymptotic notation. We write $f=O(g)$ or $f\ll g$ if there is a universal constant $C>0$ such that $|f|\leqslant Cg$. If $f$ is nonnegative and $f=O(g)$, then we also write $g=\Omega(f)$. We write $f\asymp g$ when $f\ll g$ and $g\ll f$. Finally, we write $f(p)=o(g(p))$ if $\lim_{p\to\infty}f(p)/g(p)=0$.
@@ 3. Structure theorem
Start with $E=A$. As long as $\dim(E)\geqslant R$, iteratively remove a dissociated subset of size $R/2$, so that at each step the set $E$ of remaining elements has size $|E|\geqslant R/2$. Once we reach a residual set $E$ of dimension smaller than $R$, Lemma 3.2 provides the desired $\lambda\in\mathbb{F}_{p}^{\times}$. ∎
@@
where each $D_{j}$ is a dissociated set of size $\asymp R$ and we have $E\cup\{\delta\}\subseteq(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)})$, for $\delta\vcentcolon=\sum_{j=1}^{s}\sum_{d\in D_{j}}d$. Set $P\vcentcolon=E\cap(0,p/4|E|)$ and $N\vcentcolon=E\cap(-p/4|E|,0)$. If $s=0$, then we have already obtained the desired decomposition of $\lambda\cdot A$, so for the remainder of the proof we assume that $s\geqslant 1$. Recall that $|E|\gg R$; by replacing $\lambda$ with $-\lambda$ if necessary, we may assume that $|P|\gg R$.
@@
We begin by reducing to the case where $N$ is either empty or of size at least $10$. Suppose that $100<|N|<$ 10. Note that $\sum_{n\in N}n\in(-p/90,p/90)$. Split $D_{1}$ into $|N|+1\leqslant 10$ sets each of size $\asymp R$; the remark before the proof ensures that we can absorb all of the elements of $N$ into these dissociated sets, and this procedure changes the value of $\delta$ by at most $p/90$. Notice that each newly formed $D_{j}$ still has size $\asymp R$, and that we still have $P\subseteq(0,\frac{p}{40|P\cup N|})$, $N\subseteq(-\frac{p}{40|P\cup N|},0)$, and $\delta\in(-p/40,p/40)$.
@@
Finally, suppose that $|N|\geqslant 10$, and note that also $|P|\geqslant 10$. By replacing $\lambda$ with $-\lambda$ if necessary, we may assume that $|P|\geqslant|N|$. By absorbing $5$ arbitrary elements of $N$ into $D_{1}$, we may assume that $|P|\geqslant|N|-5$. Since we still have $|N|>1$, there is some $n_{1}\in N$ such that $\delta+n_{1}\neq-\sum_{p\in P}p$; as above, we absorb $n_{1}$ into $D_{1}$, so that the final part of condition (iii) is satisfied. Now we have
@@ 4. Ordering $P$ and $N$
In [8], the first author presented a simple algorithm for inductively constructing a two-sided valid ordering of any finite subset of $\mathbb{Z}\setminus\{0\}$. This algorithm has a lot of slack, in the sense that at each step there are many possible choices. To exploit this slack and prove Proposition 4.1, we will employ a modified algorithm that greedily avoids partial sums of $\mathbf{p}$ lying in the $Y^{+}_{j}$’s and partial sums of $\mathbf{n}$ lying in the $Y^{-}_{j}$’s.
We will construct the sequences $\mathbf{p}=p_{1},\ldots,p_{|P|}$ and $\mathbf{n}=n_{1},\ldots,n_{|N|}$ from the larger indices to the smaller indices. Suppose that we have already chosen the values of $p_{|P|},p_{|P|-1},\ldots,p_{k+1}$ and $n_{|N|},n_{|N|-1},\ldots,n_{\ell+1}$. At the next step, we will choose the value of either $p_{k}$ or $n_{\ell}$. Let
@@
If we are choosing $p_{k}$, then we want $p_{k}$ not to be equal to $\pi_{k}+\delta+\nu_{\ell}$, since this choice of $p_{k}$ would lead to a zero-sum interval $p_{k-1}+\cdots+p_{1}+\delta+n_{1}+\cdots+n_{\ell}=0$. Likewise, if we are choosing $n_{\ell}$, then we want $n_{\ell}$ not to be equal to either $\delta+\nu_{\ell}$ or $\pi_{k}+\delta+\nu_{\ell}$, since these choices of $n_{\ell}$ would lead to zero-sum intervals $\delta+n_{1}+\cdots+n_{\ell-1}=0$ and $p_{k}+\cdots+p_{1}+\delta+n_{1}+\cdots+n_{\ell-1}=0$. With this in mind, we define the sets
@@
As in the algorithm in [8], the sign of the quantity $\pi_{k}+\delta+\nu_{\ell}$ will determine whether we choose the value of $p_{k}$ or the value of $n_{\ell}$ next: [^1] The apparent asymmetry in the cases arises from the assumption in Proposition 4.1 that $\delta>0$.
@@
$|\operatorname{IS}(\mathbf{p})\cap Y^{+}_{j}|\leqslant\inf_{L\in\mathbb{N}}\left(\frac{|Y^{+}_{j}|}{L}+L+4+4\sum_{i=1}^{j-1}|Y^{+}_{i}|\right)$
@@
$|\operatorname{IS}(\mathbf{n})\cap Y^{-}_{j}|\leqslant\inf_{L\in\mathbb{N}}\left(\frac{|Y^{-}_{j}|}{L}+L+4+4\sum_{i=1}^{j-1}|Y^{-}_{i}|\right)$
@@
We will prove the statement only for $|\operatorname{IS}(\mathbf{p})\cap Y^{+}_{j}|$ since the argument for $|\operatorname{IS}(\mathbf{n})\cap Y^{-}_{j}|$ is essentially identical. Recall that $\operatorname{IS}(\mathbf{p})=\{\pi_{k}=\sum_{i=1}^{k}p_{i}:0\leqslant k\leqslant|P|\}$. For $k<|P|$, write $\pi_{k}=\pi_{k+1}-p_{k+1}$. This quantity can lie in $Y^{+}_{j}$ only when the choice of $p_{k+1}$ is a $j$-step or an $i$-step for some $i<j$. We will bound these two contributions separately. Note that skip-steps and $i$-steps for $i>j$ never contribute.
@@
We first consider the contribution of $j$-steps. Notice that the partial sums $\pi_{k}$ are strictly decreasing as $k$ decreases. Suppose that the choice of $p_{k+1}$ is a $j$-step and $\pi_{k}=\pi_{k+1}-p_{k+1}\in Y^{+}_{j}$. Then we must have $\pi_{k+1}-P^{\prime}_{k+1}\subseteq Y^{+}_{j}$. Since $p_{k+1}$ is the largest element of $P^{\prime}_{k+1}$, the other $|P^{\prime}_{k+1}\setminus\{p_{k+1}\}|\geqslant k+1-3=k-2$ elements of $\pi_{k+1}-P^{\prime}_{k+1}\subseteq Y^{+}_{j}$ lie in the interval $(\pi_{k},\pi_{k+1})$; it follows that these elements are “skipped” and can never appear in $\operatorname{IS}(\mathbf{p})$. In particular, from such $j$-steps with $k>L+1$ we obtain at most $|Y^{+}_{j}|/L$ elements of $\operatorname{IS}(\mathbf{p})\cap Y^{+}_{j}$. From $j$-steps with $k\leqslant L+1$ we trivially obtain at most $L+2$ elements of $\operatorname{IS}(\mathbf{p})\cap Y^{+}_{j}$. [^2] For the argument with $\mathbf{p}$ replaced by $\mathbf{n}$, we use $L+2$ as the cutoff instead of $L+1$.
@@
We now consider the contribution of $i$-steps with $i<j$. We will trivially bound this contribution by the total number of $i$-steps with $i<j$. We claim that the number of $i$-steps is at most $3|Y_{i}|$ for each $i$. For each $i$-step $k$, let $y(k)$ denote the smallest element of $(\pi_{k}-P^{\prime}_{k})\cap Y^{+}_{i}$. It suffices to show that each $y\in Y^{+}_{i}$ appears as $y(k)$ for at most $3$ different $i$-steps $k$. If $y(k)$ is not the smallest element of $\pi_{k}-P^{\prime}_{k}$, then it is distinct from $y(k^{\prime})$ for all $k^{\prime}<k$ since
$\pi_{k^{\prime}}\leqslant\pi_{k-1}=\pi_{k}-p_{k}<y(k)$
@@
by the definition of an $i$-step. If $y(k)$ is the smallest element of $\pi_{k}-P^{\prime}_{k}$, then it is one of the two smallest elements of $\pi_{k}-P_{k}$. Notice that the smallest element of $\pi_{k}-P_{k}$ is strictly decreasing as $k$ decreases and the second-smallest element of $\pi_{k}-P_{k}$ is strictly decreasing as $k$ decreases; it follows that each $y$ can appear at most twice as one of the two smallest elements of $\pi_{k}-P_{k}$. Thus we have shown that each $i$-step $k$ in the algorithm is associated with some number $y(k)\in Y^{+}_{i}$ and moreover that any given $y\in Y^{+}_{i}$ appears as $y(k)$ for at most $3$ different $i$-steps $k$, so we conclude that the total number of $i$-steps is at most $3|Y^{+}_{i}|$. [^3] For the argument with $\mathbf{p}$ replaced by $\mathbf{n}$, the number of $i$-steps is at most $4|Y^{-}_{i}|$; we increase the constant $3$ to $4$ because $N_{\ell}\setminus N^{\prime}_{\ell}$ can now have size up to $2$, whereas $P_{k}\setminus P^{\prime}_{k}$ has size at most $1$. This establishes the claim.
Combining these contributions (and adding $1$ for $\pi_{|P|}$) gives the desired upper bound. ∎
These three claims together imply Proposition 4.1.
@@ 6. Randomizing the dissociated sets
We now choose $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ to be a uniformly random pair of orderings of $T_{2j},T_{2j+1}$ conditional on the length-$K$ prefixes of $\overline{\mathbf{t}_{2j}},\mathbf{t}_{2j+1}$ forming a permissible pair of length $(K,K)$. Equivalently, we let $t_{1},\ldots,t_{K}$ and $t^{\prime}_{1},\ldots,t^{\prime}_{K}$ be a uniformly random permissible pair of orderings of $T_{2j},T_{2j+1}$, and then we let $\mathbf{t}_{2j}$ be a uniformly random ordering of $T_{2j}$ conditional on $\overline{\mathbf{t}_{2j}}$ beginning with $t_{1},\ldots,t_{K}$, and we let $\mathbf{t}_{2j+1}$ be a uniformly random ordering of $T_{2j+1}$ conditional on $\mathbf{t}_{2j+1}$ beginning with $t^{\prime}_{1},\ldots,t^{\prime}_{K}$. We make these choices independently for different values of $j$, and independently of the choices of $\mathbf{t}_{1},\mathbf{t}_{u}$. Even though the random variables $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ are dependent, they are “conditionally anti-concentrated” in the sense that if we condition on $\mathbf{t}_{2j}$ being any particular ordering, then the the probability of $\mathbf{t}_{2j+1}$ assuming any particular ordering is still $\ll 1/m_{2j+1}!$, and vice versa. Notice that the constraints on the pairs $\mathbf{t}_{2j}\mathbf{t}_{2j+1}$ guarantee the absence of zero-sum Type I intervals beginning in the right border of $\mathbf{t}_{2j}$ and ending in the left border of $\mathbf{t}_{2j+1}$.
@@
By definition, there exists some $j$ such that $I$ contains exactly $k$ elements of $T_{j}$, where $K\leqslant k\leqslant|T_{j}|-K$. As in the first step of the proof of Lemma 5.3, we break the sum over $I$ into the sum over the part intersecting $T_{j}$ and the part not intersecting $T_{j}$ and condition on $\mathbf{t}_{i}$ for all $i\neq j$. The observations above ensure that even after this conditioning, the probability of $\mathbf{t}_{j}$ assuming any particular ordering is $\ll 1/m_{j}!$. Since the $k$-element subsets of $T_{j}$ all have distinct sums, we see that the sum over the part of $I$ intersecting $T_{j}$ assumes each particular value with probability

### v2

@@ 2. Notation and parameters
We use standard asymptotic notation. We write $f=O(g)$ or $f\ll g$ if there is a universal constant $C>0$ such that $|f|\leqslant Cg$. If $f$ is non-negative and $f=O(g)$, then we also write $g=\Omega(f)$. We write $f\asymp g$ when $f\ll g$ and $g\ll f$. Finally, we write $f(p)=o(g(p))$ if $\lim_{p\to\infty}f(p)/g(p)=0$.
@@ 3. Structure theorem
Start with $E=A$. As long as $\dim(E)\geqslant R$, iteratively remove a dissociated subset of size $R/2$, so that at each step the set $E$ of remaining elements has size $|E|\geqslant R/2$. Once we reach a residual set $E$ of dimension smaller than $R$, Lemma 3.2 (applied to $E\cup\{\delta\}$, where $\delta$ is the sum of all of the dissociated elements removed) provides the desired $\lambda\in\mathbb{F}_{p}^{\times}$. ∎
@@
where each $D_{j}$ is a dissociated set of size $\asymp R$ and we have $E\cup\{\delta\}\subseteq(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)})$, for $\delta\vcentcolon=\sum_{j=1}^{s}\sum_{d\in D_{j}}d$. Set $P\vcentcolon=E\cap(0,p/4|E|)$ and $N\vcentcolon=E\cap(-p/4|E|,0)$. If $s=0$, then we have already obtained the desired decomposition of $\lambda\cdot A$, so for the remainder of the proof we assume that $s\geqslant 1$. By replacing $\lambda$ with $-\lambda$ if necessary, we may assume that $|P|\geqslant|N|$. In particular, since $|E|\gg R$, this implies that $|P|\gg R$.
@@
We will later apply sumset inequalities involving $P,N$, and we will need $P,N$ to be not-too-small so that we have “room” for sumsets to expand. In anticipation of this, we begin by reducing to the case where $N$ is either empty or of size at least $10$. Suppose that $100<|N|<$ 10. Note that $\sum_{n\in N}n\in(-p/90,p/90)$. Split $D_{1}$ into $|N|+1\leqslant 10$ sets each of size $\asymp R$; the remark before the proof ensures that we can absorb all of the elements of $N$ into these dissociated sets, and this procedure changes the value of $\delta$ by at most $p/90$. Notice that each newly formed $D_{j}$ still has size $\asymp R$, and that we still have $P\subseteq(0,\frac{p}{40|P\cup N|})$, $N\subseteq(-\frac{p}{40|P\cup N|},0)$, and $\delta\in(-p/40,p/40)$.
@@
Finally, suppose that $|N|\geqslant 10$, and recall that we also have $|P|\geqslant|N|\geqslant 10$. By absorbing $5$ arbitrary elements of $N$ into $D_{1}$, we may assume that $|P|\geqslant|N|-5$. Since we still have $|N|>1$, there is some $n_{1}\in N$ such that $\delta+n_{1}\neq-\sum_{p\in P}p$; as above, we absorb $n_{1}$ into $D_{1}$, so that the final part of condition (iii) is satisfied. Now we have
@@ 4. Ordering $P$ and $N$
In [8], the first author presented a simple algorithm for inductively constructing a two-sided valid ordering of any finite subset of $\mathbb{Z}\setminus\{0\}$. Let us quickly review this algorithm since it forms the basis for the proof of Proposition 4.1. Suppose $P\subseteq(0,\infty)$ and $N\subseteq(-\infty,0)$ are finite sets of integers; we want to produce orderings $\mathbf{p}$ of $P$ and $\mathbf{n}$ of $N$ such that $\overline{\mathbf{p}},\mathbf{n}$ is two-sided valid. We will construct the sequences $\mathbf{p}=p_{1},\ldots,p_{|P|}$ and $\mathbf{n}=n_{1},\ldots,n_{|N|}$ from the larger indices to the smaller indices. For the first step, consider the sign of $\sum_{n\in N}n+\sum_{p\in P}p$. Suppose that this sum is non-negative; we will choose the value of $p_{|P|}$ as follows. There is some $p^{*}\in P$ such that $\sum_{n\in N}n+\sum_{p\in P\setminus\{p^{*}\}}p\neq 0$, and we choose this $p^{*}$ to be our $p_{|P|}$. This choice ensures that if $\overline{\mathbf{p}^{\prime}},\mathbf{n}$ is a two-sided valid ordering of the remaining elements $P\setminus\{p_{|P|}\},N$, then $p_{|P|},\overline{\mathbf{p}^{\prime}},\mathbf{n}$ is the desired two-sided valid ordering of $P,N$: Any interval containing $p_{|P|}$ and not containing any of $\mathbf{n}$ clearly has strictly positive sum; any proper interval containing both $p_{|P|}$ and some elements of $\mathbf{n}$ must contain all of $P$ and hence has strictly positive sum by our assumption that $\sum_{n\in N}n+\sum_{p\in P}p\geqslant 0$; the intervals strictly contained in $\overline{\mathbf{p}^{\prime}},\mathbf{n}$ are all non-zero-sum by assumption; and the interval consisting of all of $\overline{\mathbf{p}^{\prime}},\mathbf{n}$ has non-zero sum by our choice of $p^{*}$. If instead $\sum_{n\in N}n+\sum_{p\in P}p<0$, then we choose $n_{|N|}$ analogously. With this first step complete, we throw out the already-chosen element $p_{|P|}$ or $n_{|N|}$ and repeat this process with the remaining elements. This procedure produces the desired orderings $\mathbf{p},\mathbf{n}$.
The above algorithm has a lot of slack, in the sense that at each step there are many possible choices. To exploit this slack and prove Proposition 4.1, we will employ a modified algorithm that greedily avoids partial sums of $\mathbf{p}$ lying in the $Y^{+}_{j}$’s and partial sums of $\mathbf{n}$ lying in the $Y^{-}_{j}$’s.
Proof of Proposition 4.1.
We will construct the sequences $\mathbf{p}=p_{1},\ldots,p_{|P|}$ and $\mathbf{n}=n_{1},\ldots,n_{|N|}$ from the larger indices to the smaller indices. Suppose that we have already chosen the values of $p_{|P|},p_{|P|-1},\ldots,p_{k+1}$ and $n_{|N|},n_{|N|-1},\ldots,n_{\ell+1}$. At the next step, we will choose the value of either $p_{k}$ or $n_{\ell}$ depending on the sign of the sum of all of the remaining elements. Let
@@
As in the algorithm from [8], we will succeed in constructing a two-sided valid ordering as long as $p_{k},n_{\ell}$ avoid a few particular potential values (for more details, see Claim 4.3 below). If we are choosing $p_{k}$, then we want $p_{k}$ not to be equal to $\pi_{k}+\delta+\nu_{\ell}$, since this choice of $p_{k}$ would lead to a zero-sum interval $p_{k-1}+\cdots+p_{1}+\delta+n_{1}+\cdots+n_{\ell}=0$. Likewise, if we are choosing $n_{\ell}$, then we want $n_{\ell}$ not to be equal to either $\delta+\nu_{\ell}$ or $\pi_{k}+\delta+\nu_{\ell}$, since these choices of $n_{\ell}$ would lead to zero-sum intervals $\delta+n_{1}+\cdots+n_{\ell-1}=0$ and $p_{k}+\cdots+p_{1}+\delta+n_{1}+\cdots+n_{\ell-1}=0$. With this in mind, we define the sets
@@
Again as in the algorithm from [8], the sign of the quantity $\pi_{k}+\delta+\nu_{\ell}$ will determine whether we choose the value of $p_{k}$ or the value of $n_{\ell}$ next: [^1] The apparent asymmetry in the cases arises from the assumption in Proposition 4.1 that $\delta>0$.
@@
$|\operatorname{IS}(\mathbf{p})\cap Y^{+}_{j}|\leqslant\inf_{L\in\mathbb{N}}\left(\frac{|Y^{+}_{j}|}{L}+L+2+4\sum_{i=1}^{j-1}|Y^{+}_{i}|\right)$
@@
$|\operatorname{IS}(\mathbf{n})\cap Y^{-}_{j}|\leqslant\inf_{L\in\mathbb{N}}\left(\frac{|Y^{-}_{j}|}{L}+L+2+4\sum_{i=1}^{j-1}|Y^{-}_{i}|\right)$
@@
We will prove the statement only for $|\operatorname{IS}(\mathbf{n})\cap Y^{-}_{j}|$ since the argument for $|\operatorname{IS}(\mathbf{p})\cap Y^{+}_{j}|$ is essentially identical. Recall that $\operatorname{IS}(\mathbf{n})=\{\nu_{\ell}=\sum_{i=1}^{\ell}n_{i}:0\leqslant\ell\leqslant|N|\}$. For $0\leqslant\ell<|N|$, write $\nu_{\ell}=\nu_{\ell+1}-n_{\ell+1}$. This quantity can lie in $Y^{-}_{j}$ only when the choice of $n_{\ell+1}$ is a $j$-step or an $i$-step for some $i<j$. We will bound these two contributions separately. Note that skip-steps and $i$-steps for $i>j$ never contribute.
@@
We first consider the contribution of $j$-steps. Notice that the partial sums $\nu_{\ell}$ are strictly increasing (becoming less negative) as $\ell$ decreases. Suppose that the choice of $n_{\ell+1}$ is a $j$-step and $\nu_{\ell}=\nu_{\ell+1}-n_{\ell+1}\in Y^{-}_{j}$. Then we must have $\nu_{\ell+1}-N^{\prime}_{\ell+1}\subseteq Y^{-}_{j}$. Since $n_{\ell+1}$ is the smallest (most negative) element of $N^{\prime}_{\ell+1}$, the other $|N^{\prime}_{\ell+1}\setminus\{n_{\ell+1}\}|\geqslant\ell+1-3=\ell-2$ elements of $\nu_{\ell+1}-N^{\prime}_{\ell+1}\subseteq Y^{-}_{j}$ lie in the interval $(\nu_{\ell+1},\nu_{\ell})$; it follows that these elements are “skipped” and can never appear in $\operatorname{IS}(\mathbf{n})$. In particular, from such $j$-steps with $\ell\geqslant L+1$ we obtain at most $|Y^{-}_{j}|/L$ elements of $\operatorname{IS}(\mathbf{n})\cap Y^{-}_{j}$. From $j$-steps with $\ell\leqslant L$ we trivially obtain at most $L+1$ elements of $\operatorname{IS}(\mathbf{n})\cap Y^{-}_{j}$.
@@
We now consider the contribution of $i$-steps with $i<j$. We will trivially bound this contribution by the total number of $i$-steps with $i<j$. We claim that the number of $i$-steps is at most $4|Y^{-}_{i}|$ for each $i$. For each $i$-step $\ell$, let $y(\ell)$ denote the largest (least negative) element of $(\nu_{\ell}-N^{\prime}_{\ell})\cap Y^{-}_{i}$. It suffices to show that each $y\in Y^{-}_{i}$ appears as $y(\ell)$ for at most $4$ different $i$-steps $\ell$. If $y(\ell)$ is not the largest element of $\nu_{\ell}-N^{\prime}_{\ell}$, then it is distinct from $y(\ell^{\prime})$ for all $\ell^{\prime}<\ell$ since
$\nu_{\ell^{\prime}}\geqslant\nu_{\ell-1}=\nu_{\ell}-n_{\ell}>y(\ell)$
@@
by the definition of an $i$-step. If $y(\ell)$ is the largest element of $\nu_{\ell}-N^{\prime}_{\ell}$, then it is one of the three largest elements of $\nu_{\ell}-N_{\ell}$. Notice that the largest element of $\nu_{\ell}-N_{\ell}$ is strictly increasing as $\ell$ decreases, the second-largest element of $\nu_{\ell}-N_{\ell}$ is strictly increasing as $\ell$ decreases, and the third-largest element of $\nu_{\ell}-N_{\ell}$ is strictly increasing as $\ell$ decreases; it follows that each $y$ can appear at most three times as one of the three largest elements of $\nu_{\ell}-N_{\ell}$. Thus we have shown that each $i$-step $\ell$ in the algorithm is associated with some number $y(\ell)\in Y^{-}_{i}$ and moreover that any given $y\in Y^{-}_{i}$ appears as $y(\ell)$ for at most $4$ different $i$-steps $\ell$, so we conclude that the total number of $i$-steps is at most $4|Y^{-}_{i}|$. This establishes the claim.
Combining these contributions (and adding $1$ for $\nu_{|N|}$) gives the desired upper bound. [^2] To bound the intersection between $\operatorname{IS}(\mathbf{p})$ and $Y^{+}_{j}$, one simply interchanges “smaller” and “larger” throughout the proof. Since we have $|P_{k}\setminus P^{\prime}_{k}|\leq 1$ instead of $|N_{\ell}\setminus N^{\prime}_{\ell}|\leq 2$, we could replace $\ell-2$ with $k-1$ in the second paragraph and replace $4|Y_{j}^{-}|$ with $3|Y_{j}^{+}|$ in the third paragraph to obtain even a slightly tighter bound.
These three claims together imply Proposition 4.1. ∎
@@ 6. Randomizing the dissociated sets
We now choose $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ to be a uniformly random pair of orderings of $T_{2j},T_{2j+1}$ conditional on the length-$K$ prefixes of $\overline{\mathbf{t}_{2j}},\mathbf{t}_{2j+1}$ forming a permissible pair of length $(K,K)$. Equivalently, we let $t_{1},\ldots,t_{K}$ and $t^{\prime}_{1},\ldots,t^{\prime}_{K}$ be a uniformly random permissible pair of orderings of $T_{2j},T_{2j+1}$, and then we let $\mathbf{t}_{2j}$ be a uniformly random ordering of $T_{2j}$ conditional on $\overline{\mathbf{t}_{2j}}$ beginning with $t_{1},\ldots,t_{K}$, and we let $\mathbf{t}_{2j+1}$ be a uniformly random ordering of $T_{2j+1}$ conditional on $\mathbf{t}_{2j+1}$ beginning with $t^{\prime}_{1},\ldots,t^{\prime}_{K}$. We make these choices independently for different values of $j$, and independently of the choices of $\mathbf{t}_{1},\mathbf{t}_{u}$. The following lemma shows that even though the random variables $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ are dependent, they are “conditionally anti-concentrated” in the sense that if we condition on $\mathbf{t}_{2j}$ being any particular ordering, then the the probability of $\mathbf{t}_{2j+1}$ assuming any particular ordering is still very small, and vice versa.
Lemma 6.2.
Choose $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ according to the distribution described above. Then, conditional on $\mathbf{t}_{2j}$ assuming any particular ordering, the probability of $\mathbf{t}_{2j+1}$ assuming any particular ordering is $\ll 1/m_{2j+1}!$; likewise, conditional on $\mathbf{t}_{2j+1}$ assuming any particular ordering, the probability of $\mathbf{t}_{2j}$ assuming any particular ordering is $\ll 1/m_{2j}!$.
Proof.
We prove only the first statement. Let $\mathbf{u}_{2j}$ be any fixed ordering of $T_{2j}$. Let $\mathbf{u}^{(K)}_{2j}$ denote the ordering consisting of the first $K$ elements of $\mathbf{u}_{2j}$. The number of permissible pairs $\mathbf{u}^{(K)}_{2j}\mathbf{u}_{2j+1}^{(K)}$ with $\mathbf{u}_{2j+1}^{(K)}$ of length $K$ is at least
$(m_{2j+1}-K)^{K}\gg m_{2j+1}^{K},$
by our choices of $R,K$ (see the above discussion of $N(k,\ell)$). Thus, the number of orderings $\mathbf{u}_{2j+1}$ of $T_{2j+1}$ such that $\mathbf{u}^{(K)}_{2j},\mathbf{u}^{(K)}_{2j+1}$ is a permissible pair is
$\gg m_{2j+1}^{K}\cdot(m_{2j+1}-K)!\gg m_{2j+1}!.$
The lemma now follows since each of these $\gg m_{2j+1}!$ orderings of $T_{2j+1}$ is equally likely to occur as $\mathbf{t}_{2j+1}$, after we condition on $\mathbf{t}_{2j}=\mathbf{u}_{2j}$. ∎
Notice that the constraints on the pairs $\mathbf{t}_{2j}\mathbf{t}_{2j+1}$ guarantee the absence of zero-sum Type I intervals beginning in the right border of $\mathbf{t}_{2j}$ and ending in the left border of $\mathbf{t}_{2j+1}$.
@@
By definition, there exists some $j$ such that $I$ contains exactly $k$ elements of $T_{j}$, where $K\leqslant k\leqslant|T_{j}|-K$. As in the first step of the proof of Lemma 5.3, we break the sum over $I$ into the sum over the part intersecting $T_{j}$ and the part not intersecting $T_{j}$ and condition on $\mathbf{t}_{i}$ for all $i\neq j$. Lemma 6.2 ensures that even after this conditioning, the probability of $\mathbf{t}_{j}$ assuming any particular ordering is $\ll 1/m_{j}!$. Since the $k$-element subsets of $T_{j}$ all have distinct sums, we see that the sum over the part of $I$ intersecting $T_{j}$ assumes each particular value with probability